A travelling sinusoidal wave has this equation:

y(x,t) = 0.0450 m sin(25.12 m^-1x – 37.68s^-1 – 0.523)


Find the following:

A. Amplitude

B. Wave number

C. Wavelength

D. Angular frequency

E. Frequency

F. Phase angle


1
Expert's answer
2021-02-11T10:40:12-0500

A)

"A= 0.0450\ m."

B)

"k=25.12\ m^{-1}."

C)

"\lambda=\dfrac{2\pi}{k}=\dfrac{2\pi}{25.12\ m^{-1}}=0.25\ m."

D)

"\omega=37.68\ \dfrac{rad}{s}."

E)

"f=\dfrac{\omega}{2\pi}=\dfrac{37.68\ \dfrac{rad}{s}}{2\pi}=6.0\ Hz."

F)

"\phi= 0.523\ rad."